OpenStudy (anonymous):

NEED CALCULUS HELP! so I can solve c and d on a practice test using an integral and my calculator but I cannot figure out HOW to solve the actual integral by hand- which is what my teacher grades on. c) region, R, is bound by y=x^3 - 4x and y= sin(pi x). R is the base of a solid with square cross sections perpendicular to the x-axis. What is that figures volume? d) given the same R as being the surface of a pond and the knowledge that the ponds depth always = 3-x, meaning it gets more shallow as x grows and is the same depths for all values of y with a similar x. What is the pond's volume?

OpenStudy (31356):

Welcome to OpenStudy! I think it will be so much better if you had post your answer choices this way. (Example) A) 234 B) 123 C) 10,000 D) 1

OpenStudy (anonymous):

I don't have answer choices. He just told us to answer the problem. Did you maybe just mean the questions?

OpenStudy (31356):

Never mind then.......

OpenStudy (phi):

Did they give bounds for problem c) ?

OpenStudy (anonymous):

sorry, 0 to 2

OpenStudy (anonymous):

i know it is INT ( (sin (pi x) - x^3 +4x)^2) from 0 to 2 but I don't know how to solve it by hand

OpenStudy (anonymous):

I tried factoring sin (pi x) - x^3 +4x)^2 and then splitting the answer into multiple integrals so that I could do substitution but it gave me a different answer than my calculator- which is correct according to the AP site

OpenStudy (phi):

I would expand out the quadratic

OpenStudy (anonymous):

do you mean factoring (sin (pi x) - x^3 +4x)^2) so it is sin^2(pi x) - x^3 sin(pi x) + 4x sin(pi x) - x^3 sin(pi x) - x^6 -8x^4 + 4x sin(pi x) + 16x^2 ?

OpenStudy (phi):

Yes, but I don't call that factoring. we can integrate the powers of x terms easily enough the 8x sin(pi x) we can integrate by parts sin^2(pi x) we can use the identity \[ sin^2(\pi x)= \frac{1}{2} (1- \cos(2 \pi x) ) \]

OpenStudy (phi):

the hardest term is -2x^3 sin(pi x)

OpenStudy (anonymous):

for sin^2 (pi x) i had been doing int( sin^2(pi x) = - (1/pi) cos(pi x) was that wrong?

OpenStudy (phi):

yes. if we think of it as \[ u^2 du \] with u = sin(x), then du = cos(x) dx we don't have a cos(x) dx so we can't do that. we use a trig substitution

OpenStudy (anonymous):

I am sorry, but where did the pi*x go?

OpenStudy (phi):

oh, I was being sloppy. Use \[ \sin^2(\pi x)= \frac{1}{2} (1- \cos(2 \pi x) ) \]

OpenStudy (anonymous):

why is that true?

OpenStudy (phi):

cos(a+b)= cos(a)cos(b) - sin(a)sin(b) if a=b, we get cos(2a) = cos^2(a) - sin^2(a) replace cos^2(a) with 1 - sin^2(a) cos(2a) = 1 - sin^2(a)- sin^2(a) do some algebra

OpenStudy (anonymous):

thanks

OpenStudy (phi):

we can integrate x^3 sin(pi x) using integration by parts a few times in a row. Painful, but doable.

OpenStudy (anonymous):

i finally got c! thanks!

OpenStudy (phi):

mathnerd99 wrote: I'm looking back at what you've said trying to figure this out and so I have 2 questions: 1) when you said earlier sin2(πx)=1/2(1−cos(2πx)), this was in reference to 8x*sin(pi*x), this is an actual identity? also just making sure- nothing there was integrated or derived, right? The 8x sin(pi x) is not relevant here. To integrate \( \sin^2(\pi x) \) we use the identity to change it to a form we can integrate. 2)why do is an equation for cos(a+b) necessary? is it just to get to the bit about cos(2a) = 1 - sin^2(a)- sin^2(a)? would a= pi*x ? Yes. The "double angle" formula for cos(2a) follows from the sum of angles formula, and I started as far back as seemed reasonable... if you had also wanted a derivation for the sum of angles cos(a+b), I would have given you a link, because it takes too long to write out. After some work, I found the volume to be \[ 1 -\frac{24}{\pi^3} +\frac{128}{7} - \frac{256}{5} + \frac{128}{3} \approx 9.978\]